In the ballistic pendulum experiment (A.K.A conservation of momentum & energy), the velocity of the projectile was measured and recorded (vo = 5 m/s). The mass of the projectile was 12 g and the ballistic pendulum mass was 148 g. After the projectile is fired, it hits the ballistic pendulum, and both moved with velocity. Calculate the velocity LaTeX: vv of the projectile and the pendulum after the collision.
Answer #1

here,

the mass of projectile , m1 = 12 g

the mass of ballastic pendulum , m2 = 148 g

initial velocity of projectile , u = 5 m/s

using conservation of momentum

m1 * u = (m1 + m2) * v

12 * 5 = (12 + 148) * v

solving for v

v = 0.375 m/s

the velocity of pendulum and projectile after the collison is 0.375 m/s